A $75.00 Product B $64.00 C $66.00 Selling price Variable expenses: Direct materials Other variable expenses Total variable eQuestion - What is the maximum contribution margin that the company can earn per month if it makes optimal use of its 5,400 pounds of materials?
Answer #1

Calculation of maximum contribution margin

Particulars A B C
Contribution Margin per unit 30 16 19.80
Material per unit( in pounds) 22.5/3=7.5 15/3=5 12/3=4
Contribution Margin / Material 30/7.5=4 16/5=3.2 19.8/4=4.95
Ranking for manufacturing II III I

Since material is limited resource hence company will manufacture that product first which fetches maximum contribution per material.

Calculation of Maximum Contribution Margin

Product C = 900 units * 4 * 4.95 =17820

Product A = 240 units *7.5 *4 =7200

Maximum Contribution Margin = 25020

Note-----Since 5400 pounds of materials is exhausted in producing 900 units of product C (900*4=3600) and 240 units of Product A (240*7.5=1800)